### 05 Feb 2024 v1

# nnMamba: 3D Biomedical Image Segmentation, Classification and Landmark Detection with State Space Model

Abstract
In the field of biomedical image analysis, the quest for architectures capable of effectively capturing long-range dependencies is paramount, especially when dealing with 3D image segmentation, classification, and landmark detection. Traditional Convolutional Neural Networks (CNNs) struggle with locality respective field, and Transformers have a heavy computational load when applied to high-dimensional medical images. In this paper, we introduce nnMamba, a novel architecture that integrates the strengths of CNNs and the advanced long-range modeling capabilities of State Space Sequence Models (SSMs). nnMamba adds the SSMs to the convolutional residual-block to extract local features and model complex dependencies. For diffirent tasks, we build different blocks to learn the features. Extensive experiments demonstrate nnMamba’s superiority over state-of-the-art methods in a suite of challenging tasks, including 3D image segmentation, classification, and landmark detection. nnMamba emerges as a robust solution, offering both the local representation ability of CNNs and the efficient global context processing of SSMs, setting a new standard for long-range dependency modeling in medical image analysis. Code is available at https://github.com/lhaof/nnMamba
1 Introduction
Biomedical image analysis plays a pivotal role in healthcare, encompassing key tasks such as segmentation, classification, and landmark detection, which have been extensively studied in the literature [1, 2]. Image segmentation, in particular, is crucial for converting unstructured biomedical images into structured and informative representations, thereby propelling scientific discovery and clinical applications forward [3, 4]. Semantic segmentation is indispensable across various AI-enabled clinical applications, from diagnostic assistance and therapeutic planning to intra-operative support and monitoring tumor evolution [5]. Similarly, image based classification [6, 7, 8] and landmark detection tasks [9, 10] are vital for a range of downstream biomedical applications. For instance, Alzheimer’s disease (AD), a prevalent neurological disorder, can be better managed through classification models that help in understanding its progression [11]. Automated landmark detection plays a critical role in assisting radiologists by providing key reference points, thereby reducing workload and the potential for diagnostic errors [9, 12].
In the field, CNNs [13] have been a dominant force, particularly with FCN-based methods [14, 15, 16, 17] that are adept at hierarchical feature extraction. Transformers, with their origins in NLP and subsequent adaptation for visual tasks via architectures like ViT [18] and SwinTransformer [19], excel in global information aggregation. Their integration into CNN frameworks has led to hybrid models such as TransUNet [20], UNETR[21], and nnFormer [22], which significantly enhance the modeling of long-range dependencies.
However, the computational intensity of Transformers, particularly due to the self-attention mechanism’s scaling with input size, poses challenges in high-dimensional biomedical image analysis. State Space Sequence Models (SSMs), and specifically Structured State Space (S4) models [23], have emerged as efficient alternatives, delivering cutting-edge results in long-sequence data analysis [24, 23]. The Mamba model [25] refines the S4 approach by introducing an input-adaptive mechanism, outperforming Transformers in dense data scenarios. Recent works have begun exploring Mamba’s application in medical image segmentation [26, 27, 28] and visual recognition [29]. Despite these developments, a comprehensive evaluation of Mamba’s effectiveness in 3D medical segmentation, classification, and landmark detection tasks remains an open area of investigation.
Inspired by the promising results of SSMs in vision tasks and the capabilities of Mamba to handle long sequences efficiently, we explore the integration of Mamba blocks within CNNs to enhance long-range dependency modeling. Thus we proposed the nnMamba, a various of structure for both segmentation, classification, and landmark detection.
  • We present nnMamba, a framework adept at handling the intrinsic complexities of three core tasks in 3D medical imaging, namely image segmentation, classification, and landmark detection. The architecture of nnMamba is meticulously crafted to address the unique demands of each task while maintaining a cohesive structural foundation.
  • Through extensive experimentation, we demonstrate the superior performance of the nnMamba framework over existing methodologies. Our results indicate that nnMamba achieves the state-of-the-art effectiveness across the board, setting a new benchmark for future research and applications in the field of 3D medical image analysis.
2 Methodology
Figure 1: Our proposed nnMamba framework for 3D biomedical segmentation, classification and landmark detection. “BN” indicates the batch normliaztion operation.
2.1 State Space Sequence Models
State Space Sequence Models (SSMs) efficiently transform one-dimensional input sequences $u(t)$ to outputs $y(t)\in\mathbb{R}$ through a system of linear Ordinary Differential Equations (ODEs):
$x^{\prime}(t)=Ax(t)+Bu(t),$ (1)
$y(t)=Cx(t),$ (2)
with the latent state $x(t)\in\mathbb{R}^{N}$ and parameters $A\in\mathbb{R}^{N\times N}$, $B,C\in\mathbb{R}^{N}$. Despite their linear time complexity and parallelizability, SSMs often face memory and vanishing gradient issues. Structured State Space (S4) models mitigate these drawbacks by structuring the state matrix $A$ using the High-Order Polynomial Projection Operator (HIPPO), which improves long-range dependency modeling. Mamba, an extension of SSMs, introduces an input-dependent selection mechanism for adaptive parameter tuning and employs a hardware-optimized algorithm for efficient computation. This innovation leads to performance gains in processing long sequences, achieving state-of-the-art results with enhanced computational efficiency.
2.2 nnMamba for segmentation and landmark detection
In this study, we concentrate on enhancing the representational capacity of the encoder in neural network architectures through the incorporation of the Mamba layer, which is designed to capture long-range dependencies within the input data. Our methodology is applied to two distinct tasks: segmentation and landmark detection. We employ an encoder-decoder framework, utilizing a residual-based encoder coupled with a convolutional decoder. Segmentation tasks, characterized by their requirement for dense prediction, naturally demand intricate feature interactions. To address this, we introduce Res-Mamba blocks, which integrate Mamba layers subsequent to each residual block layer, thereby enhancing the network’s ability to facilitate effective feature communication at multiple scales.
For landmark detection, the objective is to generate accurate heatmaps that pinpoint anatomical landmarks. We posit that the initial stages of feature extraction are crucial for this task. Consequently, we incorporate a singular Mamba layer at the first block of the network. This strategic placement is intended to bolster the network’s proficiency in modeling long-range relationships early on in the processing pipeline, which is pivotal for the precise localization inherent to landmark detection. Our approach leverages the strengths of the Mamba layer to enrich feature representation, thereby improving performance across both segmentation and landmark detection tasks, as evidenced by our experimental results.
2.3 nnMamba for classification
In addressing the classification task, our architecture primarily employs a residual encoder. The underlying hypothesis is that when feature maps are reduced in size, the capacity to encapsulate long-range dependencies via convolutional operations is enhanced. Building on this rationale, we judiciously integrate the Mamba layer at an early stage, immediately following the initial feature extraction phase. This strategic insertion of the Mamba layer is designed to leverage the smaller spatial dimensions of the feature map, thereby enabling a more effective capture of long-range interactions. Such interactions are pivotal for classification, where the global understanding of the input data is crucial for accurate label assignment. The configuration of the Mamba layer at this juncture in the network ensures that the subsequent layers operate on features that already possess a global contextual awareness, potentially reducing the need for further complex operations to achieve high-level reasoning. Our experimental results confirm the efficacy of this approach, demonstrating enhanced classification performance attributed to the improved representational capabilities introduced by the Mamba layer.
3 Experiments
3.1 Implementation and Metric
All the models are trained with NVIDIA V100 GPU with 24GB memory. The framework is implemented in PyTorch 2.0.1[30] and CUDA 11.6. We train the models with the Adam optimizer at a learning rate of 0.002, batch size at 2, training epoch at 100, and weight decay at 0.001. For the segmentation task, we use the dice score and HD95 for evaluation. For the landmark detection task, we take the MRE as the metric for evaluation following [10]. For the classification task, we use the AUC, accuracy, recall ,specificity, precision, and F1-score by following [11].
3.2 Segmentation Evaluation
We make the evaluation of the segmentation on the BraTS 2023 GIL track [7]. We use the official training set, and split them to the training, validation, and testing with the factor of 7:1:2. We use the nnUNet framework to evaluate our code. In Table 1, we present the comparative evaluation of our proposed nnMamba network against other state-of-the-art methods on the BraTS 2023 GIL track. Performance metrics include the Dice similarity coefficient (Dice) and the 95th percentile Hausdorff Distance (HD95). Our findings demonstrate that nnMamba achieves superior performance in terms of the average Dice score, with a value of 90.01, outperforming all other methods including ViT, DIT, UNETR, and nnUNet. Specifically, nnMamba shows notable improvements in the ET category with a Dice score of 85.65. Additionally, nnMamba’s average HD95 score is competitive at 7.43. Notably, it outperforms the pioneering ViT model by a margin of 0.60 in HD95. These results underscore nnMamba’s efficacy in segmenting brain tumors with high precision, substantiating its potential as a robust tool for medical image analysis.
Table 1: Evaluation of nnMamba and other state-of-the-art methods on the BraTS 2023 GIL track.
BraTS 2023 GIL Dice HD95
Methods WT TC ET Average WT TC ET Average
ViT[18] 92.97 88.88 84.30 88.72 5.20 6.06 12.83 8.03
DIT[31] 93.49 90.22 84.38 89.36 4.21 5.27 13.64 7.71
UNETR[21] 93.33 89.89 85.19 89.47 4.76 7.27 12.78 8.27
nnUNet[5] 93.31 90.24 85.18 89.58 4.49 4.95 11.91 7.12
nnMamba 93.70 90.67 85.65 90.01 3.88 6.82 11.59 7.43
3.3 Classification Evaluation
We evaluate the classification task on the ADNI dataset[32, 33] with the AD and NC classification task. Table 2 showcases the performance of our nnMamba network compared to the 3D-ResNet baseline on the challenging ANDI (Alzheimer’s Disease Neuroimaging Initiative) classification task. The performance metrics evaluated include Accuracy, Sensitivity/Recall, Specificity, Precision, F1 score, and the Area Under the ROC Curve (AUC). The nnMamba model demonstrates a commendable improvement in Accuracy, achieving a score of 80.07, which is a relative increase from the 79.39 accuracy of the 3D-ResNet model. Remarkably, nnMamba exhibits a significant enhancement in Sensitivity/Recall, achieving an impressive 89.78 in distinguishing between Normal Control (NC) and Alzheimer’s Disease (AD) cases. This is a substantial improvement over the 77.37 sensitivity of the 3D-ResNet model, indicating the robustness of nnMamba in identifying AD cases. Specificity is lower for nnMamba (71.70) compared to 3D-ResNet (81.13), which may indicate a trade-off between identifying true positives and reducing false positives. Despite a slight decrease in Precision, nnMamba achieved a higher F1 score of 80.66, compared to 77.66 of the baseline model, suggesting a better balance between Precision and Recall. Most notably, nnMamba’s AUC of 90.04 is higher than that of 3D-ResNet, which stands at 87.90, underscoring its superior diagnostic ability in distinguishing between the two classes.
Table 2: Evaluation of nnMamba and other state-of-the-art methods on the ANDI classification task.
NC VS AD Accuracy Sensitivity/Recall Specificity Precision F1 AUC
3D-ResNet[34] 79.39 77.37 81.13 77.94 77.66 87.90
nnMamba 80.07 89.78 71.70 73.21 80.66 90.04
3.4 Landmark Detection Evaluation
For the landmark detection task, we evaluate our method on the our private fetal Brain landmark detection dataset. In Table 3, we report the performance comparison between the proposed nnMamba framework and the ResUNet baseline for anatomical landmark detection. Performance is quantified using mean error values in millimeters for landmarks. nnMamba outperforms ResUNet on HDV 1, ADV 1, and ADV 2 with mean errors of 2.23 mm, 2.05 mm, and 1.84 mm, respectively, compared to ResUNet’s 2.32 mm, 2.45 mm, and 2.05 mm in the same order. Despite nnMamba’s slightly increased mean errors in the CBD category, it shows a lower overall average mean error (2.21 mm versus ResUNet’s 2.23 mm), indicating a marginal yet consistent improvement in general landmark detection accuracy. These results underscore nnMamba’s potential as a reliable model for detecting specific anatomical landmarks, especially for HDV and ADV, corroborating its utility in medical imaging applications.
Table 3: Evaluation of nnMamba and other state-of-the-art methods on the landmark detection task.
Landmark Detection CBD 1 CBD 2 HDV 1 HDV 2 ADV 1 ADV 2 Average
ResUNet[34] 2.23 2.15 2.32 2.16 2.45 2.05 2.23
nnMamba 2.45 2.59 2.23 2.10 2.05 1.84 2.21
4 Conclusion
In summary, our work has introduced nnMamba, a novel framework that effectively fuses the granular feature extraction capabilities of CNNs with the expansive long-range dependency modeling afforded by State Space Sequence Models (SSMs). By innovatively integrating SSMs within the convolutional residual blocks, nnMamba has been tailored to excel in the domain of 3D medical image analysis, tackling image segmentation, classification, and landmark detection with unprecedented efficacy.
Our extensive experimental analysis confirms that nnMamba sets a new benchmark, surpassing existing state-of-the-art methodologies in handling complex image analysis tasks. The architecture’s capacity to seamlessly blend attention to local details with an understanding of global context makes it an essential tool in the field. The implications of nnMamba’s performance are substantial, offering a robust and versatile framework that promises to elevate the standard of medical image analysis and support the advancement of medical diagnostic and therapeutic applications.
It should be noted that current version only provides primary results. We are still working on this project by evaluating more backbones and datasets.
Acknowledgments and Disclosure of Funding
This work was supported in part by the National Natural Science Foundation of China (NO. 62102267), and in part by the Natural Science Foundation of Guangdong Province of China (2023A1515011464).